OpenStudy (anonymous):

If the length of each side of a cube is reduced to its original value, is the surface area reduced to its original value? Use an example to support your answer.

OpenStudy (anonymous):

v=l*w*h v0=l0*w0*h0 I think it would so long as each side was reduced by same constant